OpenStudy (anonymous):

For what values of h is v3 [5, -7, h] in Span {v1 =[1,-3,2] , v2 =[-3,9,-6]} and for what values of h is {v1, v2, v3} linearly dependent?

OpenStudy (anonymous):

That is to say \[\left(\begin{matrix}1 \\ -3 \\ 2\end{matrix}\right)\] = v1, and the others are vertical as well

OpenStudy (anonymous):

\[\left(\begin{matrix}\left[\begin{matrix}1 & -3 & 2 \\ 0 & 0 & 1\\ 0 & 0 & 10+h \end{matrix}\right] \\ \end{matrix}\right)\] is what I have as the REF, but I know it reduces more, just not sure how